Your division is considering two investment projects, each of which requires an up-front expenditure of $15 million. You estimate that the investments will produce the following net cash flows:

Year

Project A

Project B

1

$5,000,000

$20,000,000

2

$10,000,000

$10,000,000

3

$20,000,000

$6,000,000

a. What are the two project’s net present values, assuming the cost of capital is 5%? 10%? 15%

· 5%:

· 10%:

· 15%

b. What are the two project’s IRR at these same costs of capital?
Answer #1

at 5% cost of capital

NPVofproject A = -15000000 + 5000000/1.05+ 10000000/1.05^2 + 20000000/1.05^3 =16108951.5171




NPV of project B = -15000000 + 20000000/1.05+ 10000000/1.05^2 + 6000000/1.05^3 = 18300939.4234




at 10% cost of capital

NPVofproject A = -15000000 + 5000000/1.10+ 10000000/1.10^2 + 20000000/1.10^3 =12836213.3734



NPV of project B = -15000000 + 20000000/1.10+ 10000000/1.10^2 + 6000000/1.10^3 = 15954169.7971


at 15% cost of capital

NPVofproject A = -15000000 + 5000000/1.15+ 10000000/1.15^2 + 20000000/1.15^3 =10059587.4086



NPV of project B = -15000000 + 20000000/1.15+ 10000000/1.15^2 + 6000000/1.15^3 = 13897838.4154



IRR ofproject A = 43.97%


IRRofproject B = 82.03%